OpenStudy (cloverracer):

Solve for d. sqrtd^2-11 = 5 4, -4 5, -5 5, 6 6, -6

OpenStudy (cloverracer):

@Hero

OpenStudy (anonymous):

The equation is \[\sqrt{d^2}-11=5\] right?

OpenStudy (cloverracer):

the squareroot goes all the way to 11

OpenStudy (anonymous):

oh okay. So, you're solving \[\sqrt{d^2-11}=5\] well, the square root of d^2 is d. So, you can rewrite this as \[d-\sqrt{11}=5\] Following?

OpenStudy (cloverracer):

The squareroot of 11 is 3.31

OpenStudy (anonymous):

Right. So, what -3.31 gives you 5?

OpenStudy (cloverracer):

what do you mean what -3.31 gives you 5?

OpenStudy (anonymous):

What gives you 5 when you subtract 3.31?

OpenStudy (cloverracer):

1.69?

OpenStudy (anonymous):

Hmm.... doesn't seem to be one of your answers. Can I see the full problem? Maybe I'm not understanding this.

OpenStudy (bradely):

step by step answer is here http://www.mathskey.com/question2answer/11481/solve-for-d-sqrtd-2-11-5

OpenStudy (cloverracer):